# KarmaTS: A Universal Simulation Platform for Multivariate Time Series with Functional Causal Dynamics

### v1

@@ Abstract
We introduce KarmaTS, an interactive framework for constructing lag-indexed, executable spatiotemporal causal graphical models for multivariate time series (MTS) simulation. Motivated by the challenge of access-restricted physiological data, KarmaTS generates synthetic MTS with known causal dynamics and augments real-world datasets with expert knowledge. The system constructs a discrete-time structural causal process (DSCP) by combining expert knowledge and algorithmic proposals in a mixed-initiative, human-in-the-loop workflow. The resulting DSCP supports simulation and causal interventions, including those under user-specified distribution shifts. KarmaTS handles mixed variable types, contemporaneous and lagged edges, and modular edge functionals ranging from parameterizable templates to neural network models. Together, these features enable flexible validation and benchmarking of causal discovery algorithms through expert-informed simulation.
@@ 1.1 Motivation
Multivariate time series (MTS) arise naturally in domains as diverse as healthcare (Che et al., 2018), earth science (Angryk et al., 2020), transportation (Ghosh et al., 2009), and finance (Tsay, 2013). A hallmark of MTS is the richly interwoven network of temporal and contemporaneous interactions among variables, which causal directed graphs (DAGs), where an edge between two variables signifies a direct causal effect can be represented in an interpretable way. For instance, causal models built on electronic health records have improved early sepsis prediction (Valik et al., 2023) and characterized recovery trajectories after spinal cord injury (Ehrmann et al., 2020), while in the geosciences, they have quantified nonlinear dependencies in large observational series (Runge et al., 2019a). In engineered systems (e.g., monitoring pin voltages on a circuit board), physical connectivity directly defines the causal graph.
By contrast, most real-world applications must infer structure from data, often supplementing partial expert knowledge with causal discovery algorithms. These range from constraint-based methods (Spirtes et al., 2000) (PC, FCI) to score-based searches (GES (Chickering, 2002)) and continuous-optimization approaches (NOTEARS (Zheng et al., 2018)). However, the observed statistical distribution is often a noisy or confounded projection of the underlying physiological mechanism—shaped by latent variables, measurement error, and heterogeneous subpopulations—so purely data-driven discovery may miss dependencies that are weak in the data but structurally inevitable from anatomical or mechanistic constraints that domain experts consider unequivocal (Pearl, 2009; Runge et al., 2019a).
@@ 2.1 Synthetic MTS Generation
The CausalTime pipeline (Cheng et al., 2024b) learns a causal graph from real-world time series using a nonlinear autoregressive model to generate synthetic data. While it uses discriminative scores for fidelity, these aggregate metrics can overlook local patterns and nuances critical in expert-driven fields like healthcare.
@@ 2.2 Interactive Causal Modeling Interfaces
Several tools support visual or interactive work with causal graphs. DAGitty (Textor et al., 2016) provides a browser-based editor for static DAGs with adjustment-set analysis and bias diagnostics, and PyRCA (Liu et al., 2023) offers an interactive environment for root-cause analysis in AIOps, combining causal and dependency graphs with diagnostic workflows.
The classic Tetrad system (Scheines et al., 1994) combines a GUI for editing and visualizing graphs with a large collection of causal discovery algorithms and basic simulation tools, including time-indexed variables. However, it is organized around generic DAG/SEM and Bayesian network workflows, and time dependence is typically handled by duplicating variables across time slices rather than via a time-series–native notion of lagged and contemporaneous edges with associated functionals.
@@ 2.3 Benchmarking Platforms
CAUSEME (Runge et al., 2019a) is a web-based benchmarking platform with curated ground-truth datasets from diverse domains, featuring challenges like non-stationarity and high dimensionality. Its interface allows experts to upload data, select methods, and compare performance metrics (e.g., precision, recall), standardizing offline evaluation to rigorously assess algorithmic robustness.
@@ 2.5 Causal Discovery in Multivariate Time Series
Constraint-based methods like PCMCI (Runge et al., 2019b) and its variants (PCMCI+ (Runge, 2020), LPCMCI (Gerhardus and Runge, 2020)) use conditional independence (CI) tests to filter variables and manage confounders. Structural approaches such as VarLiNGAM (Hyvärinen et al., 2010) combine vector autoregression with LiNGAM assumptions to find effects in non-Gaussian data.
More recent scalable methods include DYNOTEARS (Pamfil et al., 2020), which extends the continuous-optimization framework of NOTEARS (Zheng et al., 2018) to time series. Deep learning approaches have also emerged: NGM (Bellot et al., 2021) uses Neural ODEs to learn causal graphs from irregularly sampled data, while TCDF (Nauta et al., 2019) applies attention-based networks to find time delays and detect confounders. Graph-neural methods like CUTS+ (Cheng et al., 2024a) employ GNNs to scale discovery for high-dimensional series.
These diverse approaches offer complementary strengths. For a comprehensive survey of the field, including datasets and evaluation metrics, see Gong et al. (Gong et al., 2024).
@@ 3.1 The Mathematical Model
We used a discrete-time structural causal process (Runge, 2020) (DSCP), defined in 3.1. All notations are described in Appendix A
@@
    $X_{t}^{j}=f^{j}\left(\operatorname{Pa}(X_{t}^{j}),\eta_{t}^{j}\right),\quad j=1,\ldots,N$
@@
  • When not specified otherwise, “parent” indicates the union of the upstream nodes from both the contemporaneous and lagged edges.
@@ A Note on the Functionals
The nature of $\bm{f}$ varies widely across different research domains, taking forms from linear to non-linear, and from simple to intricate. This diversity makes it challenging to emphasize any specific paradigm of $\bm{f}$ as universally representative, which, thankfully, is mitigated by advances in neural networks (Lu et al., 2021). DSCP, albeit complicated by the functionals, is mathematically powerful enough for modeling real-world processes (Bongers and Mooij, 2018). In Figure 5, 3 and 13, we exemplify with synthetic MTS generated by the DSCP.
@@ 3.2 Assumptions
$X_{t}^{j}=f^{j}\left(\operatorname{Pa}(X_{t}^{j})\right)+g^{j}_{t},\quad j=1,\ldots,N$
@@ 3.3 Anatomy of the Framework
The implementation of the graphical model is built upon NetworkX (Hagberg et al., 2008), which is a major Python package in network analysis.
@@
ii. Neural Network Templates
@@ 5 Example: Privacy-Conscious fMRI Synthesis
In this section, we instantiate the workflow in Figure 4 with a small, illustrative, privacy-conscious use case of fMRI data simulation. However, we do not claim any formal privacy guarantees (e.g., differential privacy Dwork et al. (2006)), and the actual privacy–utility trade-off will depend on the chosen learning objectives and fusion strategy.
@@ 5.1 Dataset and Workflow Overview
We consider a movie-watching resting-state functional Magnetic Resonance Imaging (fMRI) dataset, using the MSDL (Multi-Subject Dictionary Learning) atlas as implemented in Nilearn (Richardson et al., 2018; Abraham et al., 2014).
@@ 5.2 Instantiating the DSCP for fMRI
a. Expert Knowledge (Graphical Model)
We encode expert knowledge as a correlation-thresholded connectivity graph: nodes are brain regions and an edge is included between regions $i$ and $j$ whenever their functional connectivity $|\mathbf{C}_{ij}|$ exceeds a sparsity threshold $\tau$ following Zalesky et al. (2012). For each such pair, we add a lag-1 edge $X^{i}_{t-1}\to X^{j}_{t}$, motivated by the low temporal resolution of fMRI and fast neural dynamics, which makes a fixed single-step lag a common simplification (Smith et al., 2011; Seth et al., 2015). Self-loops with lag 1 are also included, yielding a DAG $\tilde{G}=\bigl(V,E\bigr),\;E\subseteq\{(X^{i}_{t-1},X^{j}_{t}):i,j\in[N]\}.$ For each node $X^{j}_{t}$, representing a brain region at time $t$, the parent set is $\mathrm{Pa}(X^{j}_{t})=\{X^{i}_{t-1}:(X^{i}_{t-1},X^{j}_{t})\in E\},$ as in Definition 3.1. This simple, fixed-lag construction does not identify within-lag causal ordering, but provides a concrete anatomy-informed prior for our demonstration.
b. Functionals
@@
where $\eta^{j}_{t}$ denotes stochastic innovations. In this example each $f^{j}$ is implemented as a variational autoencoder (VAE) Kingma and Welling (2013) with a gated recurrent unit (GRU) Cho et al. (2014) encoder for sequence modeling: a GRU encodes the parent trajectories into a latent distribution, from which the VAE samples during training and whose mean is used at inference, and a decoder maps latent codes back to the target outputs. Model structure and training details are given in Appendix B.1 and B.2.
@@ 5.3 Empirical Behavior and Limitations
Qualitatively, the synthetic network preserves the global organization of the real map: (i) strong left–right mirror–pair edges remain dominant over other cross-hemispheric links, consistent with homotopic interhemispheric functional connectivity (Zuo et al., 2010); (ii) the large-scale community structure resembles canonical resting-state networks reported by Yeo et al. (2011); Power et al. (2011); (iii) key long-range hub pathways align with known hub/rich-club architecture (van den Heuvel and Sporns, 2011); and (iv) although local intra-hemispheric links differ and some weights are redistributed (e.g., a somewhat stronger midline hub, slightly weaker ipsilateral columns), bilateral mirror cores remain salient. This suggests that the synthetic data can maintain key network-level properties of the original dataset while obfuscating fine-grained temporal details. We additionally report a quantitative comparison in Appendix B.3.
@@ 7.1 Overall Performance Analysis
This relative ordering is broadly consistent with established time-series causal discovery benchmarks such as CAUSEME and related studies (Runge et al., 2019a; Gong et al., 2024; Runge et al., 2019b; Pamfil et al., 2020; Nauta et al., 2019; Bellot et al., 2021; Cheng et al., 2024a): constraint-based methods like PCMCI and continuous-optimization approaches like DYNOTEARS typically achieve the strongest structural accuracy, while deep models such as TCDF and NGM exhibit more variable performance. Our KarmaTS results align with these trends at a coarse level and further clarify how performance shifts with lag, density, and latent-variable rate; Appendix F provides a brief mapping to prior benchmark reports.
@@ Acknowledgment
This research project was partially supported by the JST Moonshot R&D Program, Grant Number JPMJMS203418.
@@ B.1 VAE-based Functional Model
For a given target process $X^{j}$ with parent set $\operatorname{Pa}(X^{j})$, we model the edge functional from parent histories to the next value of $X^{j}$ using a variational autoencoder (VAE) (Kingma and Welling, 2013) with a GRU encoder.
@@ B.2 VAE Mixed Loss
$\mathcal{L}_{\mathrm{rec}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{h=1}^{H}w_{h}\,\left(\hat{y}_{b,h}-y_{b,h}\right)^{2},$
@@ B.3 Results
To quantify pattern similarity between real and synthetic connectivity, we report a “matrix correlation” (Matrix_Corr) score: the Pearson correlation between the vectorized off-diagonal entries of the two Pearson correlation matrices, following common practice in network neuroscience for comparing connectomes (Yeo et al., 2011; Power et al., 2011; van den Heuvel and Sporns, 2011). In Figure 8, this score is $0.5097$, indicating moderate alignment of the edge-strength patterns despite differences in individual weights.
@@ D.2 Simulated Datasets
  1. i
    Structural configurations. We choose three classical graph topologies—star graphs, cycle graphs, and (rooted) trees—as defined in standard graph theory (e.g., Harary (1969)). They represent distinct challenges in causal inference and reflect common patterns observed in real-world networks.
  2. $i$
    Maximum lag of edges. This parameter governs the temporal reach of causal links, allowing us to assess each algorithm’s ability to resolve delayed interactions inherent in dynamic systems. Here, we hold the edge–node ratio constant (see iii). We evaluated two lag settings: (1) small, with a maximum lag of 5 time units; and (2) large, with a maximum lag of 10 time units.
  3. iii
    Edge–node ratio (ENR). Defined as $\lvert E\rvert/\lvert V\rvert$, the ENR approximates the average number of direct causes per node. By varying the ENR (including all lagged edges), we probe how algorithms perform under different connectivity and information densities. In this experiment, the maximum lag is fixed (see ii). We consider two ENR regimes: (1) sparse, with ENR $\leq 2$; and (2) dense, with $2<\mathrm{ENR}<4$.
@@ Appendix E Factor-wise Analysis
CAUSEME (stationary, moderate lag) PCMCI, VAR-based and NOTEARS-style methods near top; deep models more variable (Runge et al., 2019a, b; Pamfil et al., 2020)
Synthetic irregular MTS (Neural ODE / NGM) NGM competitive in highly nonlinear continuous-time settings, but less stable on simpler graphs (Bellot et al., 2021)
Attention-based TCDF studies TCDF strong on some nonlinear benchmarks but not consistently dominant across all structures (Nauta et al., 2019)
High-dimensional graph benchmarks (CUTS+) CUTS+ improves scalability but may sacrifice edge-wise accuracy in dense or long-lag regimes (Cheng et al., 2024a)
@@
The result is shown in Figure 15. As the graph size increases, F1-scores decline and SID rises precipitously. PCMCI preserves the highest F1 on smaller graphs, while CUTS+ and TCDF exhibit comparatively greater robustness in SID.
.pdf
@@
Figure 16 plots F1 as the latent-variable fraction rises from $0\%$ to $80\%$. The figure shows that the effect of latent variables on performance does not follow a simple monotonic trend: both F1-score and SID fluctuate as the percentage of latent variables increases. This behavior is consistent with prior findings that latent variables do not necessarily degrade or improve causal discovery, but can either obscure or reveal structure depending on the setting and assumptions (Gerhardus and Runge, 2020; Dong et al., 2024; Cai et al., 2023).

### v2

@@ Abstract
We introduce KarmaTS, an interactive framework for constructing lag-indexed, executable spatiotemporal causal graphical models for multivariate time series (MTS) simulation. Motivated by the challenge of access-restricted physiological data, KarmaTS generates synthetic MTS with known causal dynamics and augments real-world datasets with expert knowledge. The system constructs a discrete-time structural causal process (DSCP) by combining expert knowledge and algorithmic proposals in a mixed-initiative, human-in-the-loop workflow. The resulting DSCP supports simulation and causal interventions, including those under user-specified distribution shifts. KarmaTS handles mixed variable types, contemporaneous and lagged edges, and modular edge functionals ranging from parameterization templates to neural network models. Together, these features enable flexible validation and benchmarking of causal discovery algorithms through expert-informed simulation.
@@ 1.1 Motivation
Multivariate time series (MTS) arise naturally in domains as diverse as healthcare (che2018recurrent), earth science (angryk2020multivariate), transportation (ghosh2009multivariate), and finance (tsay2013multivariate). A hallmark of MTS is the richly interwoven network of temporal and contemporaneous interactions among variables, which causal directed graphs (DAGs), where an edge between two variables signifies a direct causal effect can be represented in an interpretable way. For instance, causal models built on electronic health records have improved early sepsis prediction (valik.etal_2023) and characterized recovery trajectories after spinal cord injury (ehrmann.etal_2020), while in the geosciences, they have quantified nonlinear dependencies in large observational series (rungeInferringCausationTime2019). In engineered systems (e.g., monitoring pin voltages on a circuit board), physical connectivity directly defines the causal graph.
By contrast, most real-world applications must infer structure from data, often supplementing partial expert knowledge with causal discovery algorithms. These range from constraint-based methods (pc) (PC, FCI) to score-based searches (GES (chickering2002optimal)) and continuous-optimization approaches (NOTEARS (notears)). However, the observed statistical distribution is often a noisy or confounded projection of the underlying physiological mechanism—shaped by latent variables, measurement error, and heterogeneous subpopulations—so purely data-driven discovery may miss dependencies that are weak in the data but structurally inevitable from anatomical or mechanistic constraints that domain experts consider unequivocal (pearl2009causality; rungeInferringCausationTime2019).
@@ 2.1 Synthetic MTS Generation
The CausalTime pipeline (cheng2024causaltime) learns a causal graph from real-world time series using a nonlinear autoregressive model to generate synthetic data. While it uses discriminative scores for fidelity, these aggregate metrics can overlook local patterns and nuances critical in expert-driven fields like healthcare.
@@ 2.2 Interactive Causal Modeling Interfaces
Several tools support visual or interactive work with causal graphs. DAGitty (dagitty) provides a browser-based editor for static DAGs with adjustment-set analysis and bias diagnostics, and PyRCA (pyrca) offers an interactive environment for root-cause analysis in AIOps, combining causal and dependency graphs with diagnostic workflows.
The classic Tetrad system (tetrad) combines a GUI for editing and visualizing graphs with a large collection of causal discovery algorithms and basic simulation tools, including time-indexed variables. However, it is organized around generic DAG/SEM and Bayesian network workflows, and time dependence is typically handled by duplicating variables across time slices rather than via a time-series–native notion of lagged and contemporaneous edges with associated functionals.
@@ 2.3 Benchmarking Platforms
CAUSEME (rungeInferringCausationTime2019) is a web-based benchmarking platform with curated ground-truth datasets from diverse domains, featuring challenges like non-stationarity and high dimensionality. Its interface allows experts to upload data, select methods, and compare performance metrics (e.g., precision, recall), standardizing offline evaluation to rigorously assess algorithmic robustness.
@@ 2.5 Causal Discovery in Multivariate Time Series
Constraint-based methods like PCMCI (pcmci) and its variants (PCMCI+ (pcmci+), LPCMCI (LPCMCI)) use conditional independence (CI) tests to filter variables and manage confounders. Structural approaches such as VarLiNGAM (varlingam) combine vector autoregression with LiNGAM assumptions to find effects in non-Gaussian data.
More recent scalable methods include DYNOTEARS (dynotears), which extends the continuous-optimization framework of NOTEARS (notears) to time series. Deep learning approaches have also emerged: NGM (ngm2021) uses Neural ODEs to learn causal graphs from irregularly sampled data, while TCDF (nauta2019causal) applies attention-based networks to find time delays and detect confounders. Graph-neural methods like CUTS+ (cheng2024cuts+) employ GNNs to scale discovery for high-dimensional series.
These diverse approaches offer complementary strengths. For a comprehensive survey of the field, including datasets and evaluation metrics, see Gong et al. (gong2024causal).
@@ 3.1 The Mathematical Model
We used a discrete-time structural causal process (pcmci+) (DSCP), defined in 3.1. All notations are described in Appendix A
@@
    $X_{t}^{j}=f^{j}\left(\operatorname{Pa}(X_{t}^{j}),\eta_{t}^{j}\right),\quad j=1,\cdots,N$
@@
  • When not specified otherwise, "parents" indicates the union of the upstream nodes from both the contemporaneous and lagged edges.
@@ A Note on the Functionals
The nature of $\bm{f}$ varies widely across different research domains, taking forms from linear to non-linear, and from simple to intricate. This diversity makes it challenging to emphasize any specific paradigm of $\bm{f}$ as universally representative, which, thankfully, is mitigated by advances in neural networks (lu2021learning). DSCP, albeit complicated by the functionals, is mathematically powerful enough for modeling real-world processes (DBLP:journals/corr/abs-1803-08784). In Figure 5, 3 and 13, we exemplify with synthetic MTS generated by the DSCP.
@@ 3.2 Assumptions
$X_{t}^{j}=f^{j}\left(\operatorname{Pa}(X_{t}^{j})\right)+g^{j}_{t},\quad j=1,\cdots,N$
@@ 3.3 Anatomy of the Framework
The implementation of the graphical model is built upon NetworkX (networkx), which is a major Python package in network analysis.
@@
ii) Neural Network Templates
@@ 5 Example: Privacy-Conscious fMRI Synthesis
In this section, we instantiate the workflow in Figure 4 with a small, illustrative, privacy-conscious use case of fMRI data simulation. However, we do not claim any formal privacy guarantees (e.g., differential privacy dp_def), and the actual privacy–utility trade-off will depend on the chosen learning objectives and fusion strategy.
@@ 5.1 Dataset and Workflow Overview
We consider a movie-watching resting-state functional Magnetic Resonance Imaging (fMRI) dataset, using the MSDL (Multi-Subject Dictionary Learning) atlas as implemented in Nilearn (Richardson2018; Abraham2014).
@@ 5.2 Instantiating the DSCP for fMRI
a) Expert Knowledge (Graphical Model)
We encode expert knowledge as a correlation-thresholded connectivity graph: nodes are brain regions and an edge is included between regions $i$ and $j$ whenever their functional connectivity $|\mathbf{C}_{ij}|$ exceeds a sparsity threshold $\tau$ following Zalesky2012. For each such pair, we add a lag-1 edge $X^{i}_{t-1}\to X^{j}_{t}$, motivated by the low temporal resolution of fMRI and fast neural dynamics, which makes a fixed single-step lag a common simplification (Smith2011; Seth2015). Self-loops with lag 1 are also included, yielding a DAG $\tilde{G}=\bigl(V,E\bigr),\;E\subseteq\{(X^{i}_{t-1},X^{j}_{t}):i,j\in[N]\}.$ For each node $X^{j}_{t}$, representing a brain region at time $t$, the parent set is $\mathrm{Pa}(X^{j}_{t})=\{X^{i}_{t-1}:(X^{i}_{t-1},X^{j}_{t})\in E\},$ as in Definition 3.1. This simple, fixed-lag construction does not identify within-lag causal ordering, but provides a concrete anatomy-informed prior for our demonstration.
b) Functionals
@@
where $\eta^{j}_{t}$ denotes stochastic innovations. In this example each $f^{j}$ is implemented as a variational autoencoder (VAE) kingma2013auto with a gated recurrent unit (GRU) cho2014learning encoder for sequence modeling: a GRU encodes the parent trajectories into a latent distribution, from which the VAE samples during training and whose mean is used at inference, and a decoder maps latent codes back to the target outputs. Model structure and training details are given in Appendix B.1 and B.2.
@@ 5.3 Empirical Behavior and Limitations
Qualitatively, the synthetic network preserves the global organization of the real map: (i) strong left–right mirror–pair edges remain dominant over other cross-hemispheric links, consistent with homotopic interhemispheric functional connectivity (Zuo2010); (ii) the large-scale community structure resembles canonical resting-state networks reported by Yeo2011; Power2011; (iii) key long-range hub pathways align with known hub/rich-club architecture (vandenHeuvel2011); and (iv) although local intra-hemispheric links differ and some weights are redistributed (e.g., a somewhat stronger midline hub, slightly weaker ipsilateral columns), bilateral mirror cores remain salient. This suggests that the synthetic data can maintain key network-level properties of the original dataset while obfuscating fine-grained temporal details. We additionally report a quantitative comparison in Appendix B.3.
@@ 7.1 Overall Performance Analysis
This relative ordering is broadly consistent with established time-series causal discovery benchmarks such as CAUSEME and related studies (rungeInferringCausationTime2019; gong2024causal; pcmci; dynotears; nauta2019causal; ngm2021; cheng2024cuts+): constraint-based methods like PCMCI and continuous-optimization approaches like DYNOTEARS typically achieve the strongest structural accuracy, while deep models such as TCDF and NGM exhibit more variable performance. Our KarmaTS results align with these trends at a coarse level and further clarify how performance shifts with lag, density, and latent-variable rate; Appendix F provides a brief mapping to prior benchmark reports.
@@ Acknowledgment
This research project was partially supported by the JST Moonshot R&D Program, Grant Number JPMJMS2034-18.
@@ B.1 VAE-based Functional Model
For a given target process $X^{j}$ with parent set $\operatorname{Pa}(X^{j})$, we model the edge functional from parent histories to the next value of $X^{j}$ using a variational autoencoder (VAE) (kingma2013auto) with a GRU encoder.
@@ B.2 VAE Mixed Loss
$\mathcal{L}_{\mathrm{rec}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{h=1}^{H}w_{h}\,\bigl(\hat{y}_{b,h}-y_{b,h}\bigr)^{2},$
@@ B.3 Results
To quantify pattern similarity between real and synthetic connectivity, we report a “matrix correlation” (Matrix_Corr) score: the Pearson correlation between the vectorized off-diagonal entries of the two Pearson correlation matrices, following common practice in network neuroscience for comparing connectomes (Yeo2011; Power2011; vandenHeuvel2011). In Figure 8, this score is $0.5097$, indicating moderate alignment of the edge-strength patterns despite differences in individual weights.
@@ D.2 Simulated Datasets
  1. i)
    Structural configurations. We choose three classical graph topologies—star graphs, cycle graphs, and (rooted) trees—as defined in standard graph theory (e.g. harary1969graph). They represent distinct challenges in causal inference and reflect common patterns observed in real-world networks.
  2. ii)
    Maximum lag of edges. This parameter governs the temporal reach of causal links, allowing us to assess each algorithm’s ability to resolve delayed interactions inherent in dynamic systems. Here, we hold the edge–node ratio constant (see iii)). We evaluated two lag settings: (1) small, with a maximum lag of 5 time units; and (2) large, with a maximum lag of 10 time units.
  3. iii)
    Edge–node ratio (ENR). Defined as $\lvert E\rvert/\lvert V\rvert$, the ENR approximates the average number of direct causes per node. By varying the ENR (including all lagged edges), we probe how algorithms perform under different connectivity and information densities. In this experiment, the maximum lag is fixed (see ii)). We consider two ENR regimes: (1) sparse, with ENR $\leq 2$; and (2) dense, with $2<\mathrm{ENR}<4$.
@@ Appendix E Factor-wise Analysis
CAUSEME (stationary, moderate lag) PCMCI, VAR-based and NOTEARS-style methods near top; deep models more variable (rungeInferringCausationTime2019; pcmci; dynotears)
Synthetic irregular MTS (Neural ODE / NGM) NGM competitive in highly nonlinear continuous-time settings, but less stable on simpler graphs (ngm2021)
Attention-based TCDF studies TCDF strong on some nonlinear benchmarks but not consistently dominant across all structures (nauta2019causal)
High-dimensional graph benchmarks (CUTS+) CUTS+ improves scalability but may sacrifice edge-wise accuracy in dense or long-lag regimes (cheng2024cuts+)
@@
The result is shown in Figure 15. As the graph size increases, F1 Scores decline and SID rises precipitously. PCMCI preserves the highest F1 on smaller graphs, while CUTS+ and TCDF exhibit comparatively greater robustness in SID.
@@
Figure 16 plots F1 as the latent-variable fraction rises from $0\%$ to $80\%$. The figure shows that the effect of latent variables on performance does not follow a simple monotonic trend: both F1-score and SID fluctuate as the percentage of latent variables increases. This behavior is consistent with prior findings that latent variables do not necessarily degrade or improve causal discovery, but can either obscure or reveal structure depending on the setting and assumptions (LPCMCI; dong2024rlcd; cai2023cumulants).